Help me solve to calculate the unit costs! if you could help with other requirements too!Tribute Inc. makes two styles of trophies, basic and deluxe, and operates at capacity. Tribute does large custom orders. Trib

Tribute Inc. makes two styles of​ trophies, basic and​ deluxe, and operates at capacity.Tribute does large custom orders. Tributebudgets to produce 10,000basic trophies and 5,000 deluxe trophies. Manufacturing takes place in two production​ departments: forming and assembly. In the forming​ department, indirect manufacturing costs are accumulated in two cost​ pools, set up and general overhead. In the assembly​ department, all indirect manufacturing costs are accumulated in one general overhead cost pool. The basic trophies are formed in batches of 200 but because of the more intricate detail of the deluxe​ trophies, they are formed in batches of 50.
Answer #1
The parts solved already are all correct
Remaining are as follows:
Calculation of units cost
Basic Deluxe Total
Total Direct Costs           80,700           66,500        147,200
Allocated Overhead
Forming           35,168           20,272           55,440
Assembly           21,560           26,040           47,600
Total costs        137,428        112,812        250,240
Unit cost 13.7428 22.5624
3.Activity rate = Overhead/Total Activity Rate
Set up costs 15840 150 105.6 per Set up
General overhead forming 39600 49500 0.8 per DL$
General overhead Assembly 47600 34000 1.4 per DL$
Calculation of unit cost
Basic Deluxe Total
Total Direct Costs           80,700           66,500        147,200
Allocated Overhead
Set up costs             5,280           10,560           15,840
General overhead forming           25,120           14,480           39,600
General overhead Assembly           21,560           26,040           47,600
Total costs        132,660        117,580        250,240
Unit cost 13.266 23.516

Since different rates are used to allocate overheads, costs are different among various methods

ABC System is the most accurate as it better represents cost consumption

5 Total Basic 80700 Deluxe 66500 =B6+C6 =1.12*31400 =1.4*15400 =SUM(B6:39) =B10/10000 =1.12*18100 =1.4*18600 =SUM(C6:09) =C10